# Twists of Superconformal Algebras

### v1

@@ 4.6. Case Study: 4d $\mathcal{N}=2$ Chiral Supercharges
$S_{+}\otimes W_{2}=\mathrm{Span}(\mathcal{Q}_{\alpha\mathcal{I}}),\quad S_{-}\otimes W_{2}=\mathrm{Span}(\widetilde{\mathcal{S}}^{\dot{\alpha}\mathcal{I}})$
@@
where $\alpha\in\{+,-\}$, $\dot{\alpha}\in\{\dot{+},\dot{-}\}$ and $\mathcal{I}\in\{1,2\}$. Rank $(2,0)$ elements contained in the first summand $S_{+}\otimes W_{2}$ lie in an ordinary 4d $\mathcal{N}=2$ super Poincaré algebra; this has the nature of the Donaldson–Witten twisting supercharge in [Witten_1988]. Beem et al [Beem_2015] refer to rank $(2,0)$ elements not lying in either summand in terms of “Schur” supercharges, whose name comes from their connection to the Schur limit of the superconformal index of 4d $\mathcal{N}=2$ superconformal theories, as Beem et al explain in the introduction and Section 4.4 of loc. cit. Following their terminology, a “Schur operator” in a superconformal theory is a local operator representing a non-trivial cohomology class for a fixed Schur supercharge.
4.6.1. Relationship with Donaldson–Witten supercharges
@@
As we have seen, all rank $(2,0)$ supercharges $\mathcal{Q}$ lie in the same orbit for the complex group $\mathrm{SO}(6,\mathbb{C})\times\mathrm{Sp}(2,\mathbb{C})$, and the complex Lie algebras $\mathfrak{z}^{\mathbb{C}}_{\mathcal{Q}},\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$ are conjugate for different choices of $\mathcal{Q}$. From the complexified point of view there is no distinction between twisting by a Donaldson–Witten supercharge and a Schur supercharge for any specific affine patch. This is not true at the level of a real form (i.e. for the actual group acting on the conformal compactification space in any specified signature), where there is an interesting orbit space as we saw in the preceding subsection.
@@
Indeed, if we choose a specific affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$, and therefore a subalgebra $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$, as described above, the Schur supercharge does not lie in the same $\mathrm{Spin}(p,4-p)$ orbit as the Donaldson–Witten supercharges. It is worth remarking, however, that this distinction relied on fixing an affine patch. On the other hand, given any rank $(2,0)$ supercharge, we may always choose an affine patch for which it has Donaldson–Witten type.
@@
In other words, such a Schur supercharge on one “spacetime” (i.e., on one affine patch) looks like a Donaldson–Witten supercharge after moving to another “spacetime” by a conformal transformation.
@@
Consider now the image of $[\mathcal{Q},-]$ inside the even part of the complexified superconformal algebra. It is isomorphic, as a vector space, to $\mathbb{C}^{4}$. In fact, it is an abelian Lie subalgebra, and we are free to view it as the complexified translations of an affine slice, so let us denote it by $T_{\mathbb{C}}$. (There is another copy of $\mathbb{C}^{4}$, orthogonal to this image, which we can then view as the special conformal transformations for this affine slice.)
Even better, as $\mathcal{Q}$ comes from a real form, it is a fixed point of the corresponding involution (i.e., of the appropriate version of complex conjugation). The bracket is equivariant for this involution, so the image of $[\mathcal{Q},-]$ restricted to the odd part of the real form gives a 4-dimensional real subspace $T_{\mathbb{R}}$ of $T_{\mathbb{C}}$. We define an affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$ as the inclusion of the open orbit of the $T_{\mathbb{R}}$-action on the conformal compactification $C(\mathbb{R}^{p,4-p})$. ∎
4.6.2. Real Forms of the Schur supercharges
Let us now discuss the behavior of a Schur supercharge in various real forms, along with the real groups $Z_{\mathcal{Q}}$ and $B_{\mathcal{Q}}$. For concreteness, let us work in the basis described above and consider the supercharge
$\mathbb{Q}=\mathcal{Q}_{+1}+\mathcal{S}^{\dot{+}2}.$
Alternatively, viewing this supercharge as a $2\times 4$ matrix via the given bases for $S_{+},W$, we can identify
$\mathbb{Q}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\end{pmatrix}.$
In this basis we can identify the complex Lie algebras of $\mathbb{Q}$-closed and $\mathbb{Q}$-exact elements explicitly. We now describe the projections of these algebras onto the Lie algebra of complexified infinitesimal conformal transformations:
$\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{h}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
$\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{k}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
with $\mathfrak{k}\subseteq\mathfrak{h}\subseteq\mathfrak{gl}(2,\mathbb{C})$ where $\mathfrak{k}$ denotes the subalgebra consisting of matrices with non-zero elements in the left column only and where $\mathfrak{h}$ denotes the subalgebra of lower triangular matrices.
4.6.3. Signature $(4,0)$
$\theta_{\mathrm{Euc}}(X)=-JXJ$
@@ 4.6.3. Signature $(4,0)$
The fixed point locus of $\theta_{\mathrm{Euc}}$ acting on $\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ is three complex-dimensional, or six real-dimensional. It consists of matrices of the form
$\mathfrak{z}_{\mathbb{Q}}\cap\mathfrak{so}(1,5)=\left\{\left(\begin{array}[]{c|c}a\mathbb{I}&b\mathbb{I}\\ \hline\cr c\mathbb{I}&d\mathbb{I}\end{array}\right)\colon a,b,c,d\in\mathbb{C},a+d=0\right\}$
The fixed point locus of $\theta_{\mathrm{Euc}}$ acting on $\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ is trivial, so there are no real $\mathbb{Q}$-exact conformal transformations in Euclidean signature.
4.6.4. Signature $(3,1)$
$\theta_{\mathrm{Lor}}(X)=-KX^{\dagger}K$
where
$K=\mathrm{diag}(1,1,-1,-1).$
The fixed point locus of $\theta_{\mathrm{Lor}}$ acting on $\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ is seven real-dimensional. It consists of matrices of the form
$\mathfrak{z}_{\mathbb{Q}}\cap\mathfrak{so}(2,4)$ =$=\Bigg{\{}\left(\begin{array}[]{c|c}A&B\\ \hline\cr\overline{B}&D\end{array}\right)\colon A=\mathrm{diag}(\alpha,\beta),D=\mathrm{diag}(\gamma,\delta),B=\mathrm{diag}(a,b),$
This subalgebra of $\mathfrak{su}(2,2)$ is a semisimple quotient isomorphic to $\mathfrak{su}(2)\oplus\mathfrak{su}(2)$, and the solvable radical generates a dilation in the conformal group. The two factors of this semisimple quotient act on the two factors of $\mathbb{C}\oplus\mathbb{C}\cong\mathbb{R}^{4}$ by antiholomorphic vector fields.
The fixed point locus of $\theta_{\mathrm{Lor}}$ acting on $\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ is three real-dimensional. It consists of matrices of the form
This Lie algebra is isomorphic to $\mathfrak{su}(2)$ (embedded in the $\mathfrak{su}(2)\oplus\mathfrak{su}(2)$ above via the map $X\mapsto(X,0)$), and therefore acts on an embedded complex line $\mathbb{C}\subseteq\mathbb{C}^{2}\cong\mathbb{R}^{4}$ by antiholomorphic vector fields.
Remark 4.26.
4.6.5. Signature $(2,2)$
We will only discuss the split signature case very briefly. Recall that we must consider the real form $\mathfrak{sl}(4,\mathbb{R})\subseteq\mathfrak{sl}(4,\mathbb{C})$, so the intersections $\mathfrak{z}_{\mathbb{Q}}\cap\mathfrak{so}(3,3)$ and $\mathfrak{b}_{\mathbb{Q}}\cap\mathfrak{so}(3,3)$ are very easy to understand: they are the subalgebras of $\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ and $\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}}\cap\mathfrak{sl}(4,\mathbb{C})$ consisting of matrices with real coefficients. In particular, if we identify the upper-right block matrices with translations, then there are three closed real translations, of which two are exact.

### v2

@@ 5.3. The Case of 4d $\mathcal{N}=2$ Chiral Supercharges
$S_{+}\otimes W_{2}=\mathrm{Span}(\mathcal{Q}_{\alpha\mathcal{I}})\quad\text{and}\quad S_{-}\otimes W_{2}=\mathrm{Span}(\widetilde{\mathcal{S}}^{\dot{\alpha}\mathcal{I}})$
where $\alpha\in\{+,-\}$, $\dot{\alpha}\in\{\dot{+},\dot{-}\}$, and $\mathcal{I}\in\{1,2\}$.
For example, in this basis, the supercharge $\mathcal{Q}_{+1}$ is a Donaldson–Witten supercharge [Witten_1988]. Note that it lies in the first summand $S_{+}\otimes W_{2}$, and hence in an ordinary 4d $\mathcal{N}=2$ super Poincaré algebra. Thus it acts on supersymmetric theories, not just superconformal theories, and famously twists $\mathcal{N}=2$ super Yang–Mills theory into a 4d topological field theory.
As another example, the supercharge
$\mathbb{Q}=\mathcal{Q}_{+1}+\widetilde{\mathcal{S}}^{\dot{+}2}$
provides a Schur supercharge, which we will examine in more detail below.
@@
Beem et al [Beem_2015] refer to any rank $(2,0)$ elements the do not lie in either summand as Schur supercharges. The name comes from their connection to the Schur limit of the superconformal index of 4d $\mathcal{N}=2$ superconformal theories. (See the introduction and Section 4.4 of loc. cit. for explanation of this context.) Following their terminology, a “Schur operator” in a superconformal theory is a local operator representing a non-trivial cohomology class for a fixed Schur supercharge.
Beem al discovered that these Schur operators yield vertex algebras, providing a beautiful and powerful relationship between 4d and 2d field theory. We return to this aspect in Section 8, where we explain how conformal twists modify observables of superconformal field theories. In particular, we discuss in Section 8.4 some powerful results of Saberi and Williams on this topic that provide many examples of how factorization techniques and twisting combine to yield vertex algebras, which are complementary to the structural work we do in Section 8.
In this section we will analyze two separate problems:
  • For the Schur twist, how does the signature interact with the cohomology? In particular, which real conformal transformations are exact and which are closed?
  • Is there a best affine patch for any given supercharge? We examine whether changes of affine patch allow one to recognize a twist as a Donaldson–Witten supercharge.
Analogs of these questions are relevant for other superconformal algebras too.
5.3.1. The real forms of the cohomology of a Schur supercharge
Let us now discuss how a Schur supercharge interacts with the various real forms, along with the real groups $Z_{\mathcal{Q}}$ and $B_{\mathcal{Q}}$. For concreteness, let us work in the basis described above and consider the supercharge
$\mathbb{Q}=\mathcal{Q}_{+1}+\widetilde{\mathcal{S}}^{\dot{+}2}.$
Alternatively, viewing this supercharge as a $2\times 4$ matrix via the given bases for $S_{\pm},W_{2}$, we can identify
(2) $\mathbb{Q}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\end{pmatrix}$
where the first $2\times 2$ block relates to $S_{+}$ and the second block to $S_{-}$. In this basis we can identify the complex Lie algebras of $\mathbb{Q}$-closed and $\mathbb{Q}$-exact elements explicitly.
We now describe the projections of these algebras onto the Lie algebra of complexified infinitesimal conformal transformations $\mathfrak{sl}(4,\mathbb{C})$, as we are particularly interested in how the Schur twist interacts with conformal geometry.
Let $\pi:\mathfrak{sconf}(4|2,\mathbb{C})_{0}\to\mathfrak{sl}(4,\mathbb{C})$ denote projection onto the conformal Lie algebra, i.e., forget the R-symmetries. We will view elements of $\mathfrak{sl}(4,\mathbb{C})$ as block matrices whose blocks are $2\times 2$ matrices.
Lemma 5.14.
The projection of the closed and exact subalgebras are
$\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{p}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
$\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{k}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
where
$\mathfrak{k}\subseteq\mathfrak{p}\subseteq\mathfrak{gl}(2,\mathbb{C})$
with
  • $\mathfrak{p}$ the subalgebra of lower triangular matrices, and
  • $\mathfrak{k}$ the subalgebra of matrices with non-zero elements in the left column only.
This lemma identifies the cohomology, with the $B$ and $C$ factors sitting in $\mathfrak{p}/\mathfrak{k}$, and with the $A$ and $D$ factors living in a quotient more involved to describe, due to the trace condition.
Signature $(4,0)$
$\theta_{\mathrm{Euc}}(X)=-JXJ$
where
The fixed point locus of $\theta_{\mathrm{Euc}}$ acting on $\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})$ is three complex-dimensional, or six real-dimensional. It consists of matrices of the form
$\pi(\mathfrak{z}_{\mathbb{Q}})\cap\mathfrak{so}(1,5)=\left\{\left(\begin{array}[]{c|c}a\mathbb{I}&b\mathbb{I}\\ \hline\cr c\mathbb{I}&d\mathbb{I}\end{array}\right)\colon a,b,c,d\in\mathbb{C},a+d=0\right\}$
The fixed point locus of $\theta_{\mathrm{Euc}}$ acting on $\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})$ is the zero matrix, so there are no nonzero real $\mathbb{Q}$-exact conformal transformations in Euclidean signature.
In particular, the cohomology is precisely $\pi(\mathfrak{z}_{\mathbb{Q}})\cap\mathfrak{so}(1,5)$. This result tells us how the twists interact with the conformal geometry.
Signature $(3,1)$
$\theta_{\mathrm{Lor}}(X)=-KX^{\dagger}K$
where
$K=\mathrm{diag}(1,1,-1,-1).$
The fixed point locus of $\theta_{\mathrm{Lor}}$ acting on $\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})\cap\mathfrak{sl}(4,\mathbb{C})$ is seven real-dimensional. It consists of matrices of the form
$\pi(\mathfrak{z}_{\mathbb{Q}})=\Bigg{\{}\left(\begin{array}[]{c|c}A&B\\ \hline\cr\overline{B}&D\end{array}\right)$ $\colon A=\mathrm{diag}(\alpha,\beta),D=\mathrm{diag}(\gamma,\delta),B=\mathrm{diag}(a,b),$
This subalgebra of $\mathfrak{su}(2,2)$ has a semisimple quotient isomorphic to $\mathfrak{su}(2)\oplus\mathfrak{su}(2)$, and the solvable radical generates a dilation in the conformal group. The two factors of this semisimple quotient act on the two factors of $\mathbb{C}\oplus\mathbb{C}\cong\mathbb{R}^{4}$ by antiholomorphic vector fields.
The fixed point locus of $\theta_{\mathrm{Lor}}$ acting on $\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})$ is three real-dimensional. It consists of matrices of the form
This Lie algebra is isomorphic to $\mathfrak{su}(2)$, viewed as a subalgebra of $\mathfrak{su}(2)\oplus\mathfrak{su}(2)$ via the map $X\mapsto(X,0)$, and therefore acts on an embedded complex line $\mathbb{C}\subseteq\mathbb{C}^{2}\cong\mathbb{R}^{4}$ by antiholomorphic vector fields.
Remark 5.15.
Signature $(2,2)$
We will only discuss the split signature case very briefly. Recall that we must consider the real form $\mathfrak{sl}(4,\mathbb{R})\subseteq\mathfrak{sl}(4,\mathbb{C})$, so the projections $\pi(\mathfrak{z}_{\mathbb{Q}})$ and $\pi(\mathfrak{b}_{\mathbb{Q}})\cap\mathfrak{so}(3,3)$ are very easy to understand: they are the subalgebras of $\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})\cap\mathfrak{sl}(4,\mathbb{C})$ and $\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})\cap\mathfrak{sl}(4,\mathbb{C})$ consisting of matrices with real coefficients. In particular, if we identify the upper-right block matrices with translations, then there are three closed real translations, of which two are exact.
5.3.2. Relationship with Donaldson–Witten supercharges and real twists
@@
As we have seen, all rank $(2,0)$ supercharges $\mathcal{Q}$ lie in the same orbit for the complex group $\mathrm{SO}(6,\mathbb{C})\times\mathrm{Sp}(2,\mathbb{C})$, and the complex Lie algebras $\mathfrak{z}^{\mathbb{C}}_{\mathcal{Q}}$ and $\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$ are conjugate for different choices of $\mathcal{Q}$. Hence, from this complexified point of view, and on any specific affine path, there is no distinction between twisting by a Donaldson–Witten supercharge and a Schur supercharge: we can act by the complex superconformal group to relate them.
@@
There is a distinction at the level of a real form of the superconformal group (i.e. for the actual conformal group acting on the conformal compactification space in any specified signature). As we saw in the preceding subsection, supercharges can live in different orbits for such a real group. Indeed, if we choose a specific affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$, and therefore a subalgebra $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$, as described above, the Schur supercharge does not lie in the same $\mathrm{Spin}(p,4-p)$ orbit as the Donaldson–Witten supercharges. It is worth remarking, however, that this distinction relied on fixing an affine patch.
On the other hand, given a real rank $(2,0)$ supercharge, we may always choose an affine patch for which it has Donaldson–Witten type.
@@
In other words, any real supercharge on one “spacetime” (i.e., on one affine patch) looks like a Donaldson–Witten supercharge after moving to another “spacetime” by a conformal transformation.
@@
Consider now the image $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$ of $[\mathcal{Q},-]$ inside the even part of the complexified superconformal algebra. It is isomorphic, as a Lie algebra, to $\mathfrak{sl}(2,\mathbb{C})\ltimes(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, and we are free to view the four dimensional abelian factor as the complexified translations of an affine slice, so let us denote this subalgebra by $T_{\mathbb{C}}$. (There is another copy of $\mathbb{C}^{4}$, orthogonal to this image, which we can then view as the special conformal transformations for this affine slice.)
Even better, as $\mathcal{Q}$ comes from a real form, it is a fixed point of the corresponding involution (i.e., of the appropriate version of complex conjugation). The bracket is equivariant for this involution, so the image of $[\mathcal{Q},-]$ restricted to the odd part of the real form gives us a 4-dimensional real subspace $T_{\mathbb{R}}$ of $T_{\mathbb{C}}$. We define an affine patch $\iota\colon\mathbb{R}^{p,4-p}\hookrightarrow C(\mathbb{R}^{p,4-p})$ as the inclusion of the open orbit of the $T_{\mathbb{R}}$-action on the conformal compactification $C(\mathbb{R}^{p,4-p})$. ∎
This result may appear quite surprising, as it seems to suggest that one can convert a Schur supercharge to a Donaldson-Witten supercharge by picking a different affine patch (i.e., linear spacetime). But, in fact, no Schur charge is real in Lorentzian or Euclidean signature. Thus the proposition does not apply to Schur-twisted theories.
Lemma 5.17.
The only twist in the real super Lie algebra $\mathfrak{sconf}(3,1|2)$ is the zero element.
Proof.
In Lorentzian signature, we have $\mathfrak{sconf}(3,1|2)=\mathfrak{su}(2,2|2)$ where
$\mathfrak{su}(2,2|2)=\{A\in\mathfrak{gl}(4|2,\mathbb{C})\,:\,A\mathbb{K}+(-1)^{|A|}\mathbb{K}A=0,\,\mathrm{str}(A)=0\}$
@@
$\mathbb{K}={\rm diag}(1,1,-1,-1,1,1),$
where the first four columns are even and the last two columns are odd. Let $K={\rm diag}(1,1,-1,-1)$ denote the signature $(2,2)$ form (or matrix) acting on the even vector space.
An odd element $\mathbb{Q}\in\mathfrak{su}(2,2|2)$ thus has the block form
$\left(\begin{array}[]{c|c}0&KQ^{\dagger}\\ \hline\cr QK&0\end{array}\right)$
with $Q\in\mathrm{Mat}_{2,2}(\mathbb{C})$ and $Q^{\dagger}$ its conjugate-transpose. Then
$[\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}KQ^{\dagger}QK&0\\ \hline\cr 0&QQ^{\dagger}\end{array}\right)$
by direct computation. Note that the diagonal entries of $QQ^{\dagger}$ are the norms of each row of $Q$, as a vector in $\mathbb{C}^{2}$ with the standard sesquilinear pairing. In other words, these diagonal entries are the norms as vectors in Euclidean $\mathbb{R}^{4}$ and hence must be nonnegative. We can then have $[\mathbb{Q},\mathbb{Q}]=0$ if and only if every row is zero, i.e., $Q=0$. ∎
On the other hand, in Euclidean signature, there is a space of real twists. Recall that $\mathfrak{sconf}(4,0|2)=\mathfrak{sl}(2|1,\mathbb{H})$.
Lemma 5.18.
A twist in $\mathfrak{sconf}(4,0|2)$ has the form
$\left(\begin{array}[]{c|c}0&0\\ \hline\cr C&0\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{c|c}0&B\\ \hline\cr 0&0\end{array}\right)$
in $\mathfrak{sl}(2|1,\mathbb{H})$.
This result lets us show the following.
Corollary 5.19.
The Schur supercharge does not lie in $\mathfrak{sl}(2|1,\mathbb{H})\subset\mathfrak{sl}(4|2,\mathbb{C})$.
Proof of corollary.
To see this, we merely need to identify which complex matrices appear in the embedding. Observe that the odd component of $\mathfrak{sl}(4|2,\mathbb{C})$ is isomorphic to
$\hom_{\mathbb{C}}(\mathbb{C}^{2},\mathbb{C}^{4})\oplus\hom_{\mathbb{C}}(\mathbb{C}^{4},\mathbb{C}^{2})$
and we view the Schur supercharge as living in the second summand. Recall the matrix (2) given above.
We need to describe how quaternionic matrices sit inside complex matrices. In particular, we want to describe
$\hom_{\mathbb{H}}(\mathbb{H}^{2},\mathbb{H})\hookrightarrow\hom_{\mathbb{C}}(\mathbb{C}^{4},\mathbb{C}^{2})$
explicitly. As a first step, identify $\mathbb{C}^{2}$ with $\mathbb{H}$ by sending a column vector $(a,b)^{T}$ of complex numbers to $a+jb$ (note that $j$ precedes $b$). Then an $\mathbb{H}$-linear map $f:\mathbb{H}^{2}\to\mathbb{H}$ is determined by a $2\times 1$ matrix
$(a+jb\;c+jd)$
for some complex numbers $a,b,c,d$. Thus we have an inclusion
$\hom_{\mathbb{H}}(\mathbb{H}^{2},\mathbb{H})\hookrightarrow\hom_{\mathbb{C}}(\mathbb{C}^{2},\mathbb{C}^{4})$
sending $f$ where $f(1)=(a+jb,c+jd)$ to the matrix
$\begin{pmatrix}a&b&c&d\\ -b&a&-d&c\end{pmatrix}.$
The Schur matrix (2) manifestly does not have this form. (By contrast, the usual Donaldson–Witten twist has $a=1$ and all other entries zero.) ∎
Proof of lemma.
An odd element has the block form
$\mathbb{Q}=\left(\begin{array}[]{c|c}0&B\\ \hline\cr C&0\end{array}\right)$
where $B$ is any $2\times 1$-matrix and $C$ is any $1\times 2$-matrix. Then
$[\mathbb{Q},\mathbb{Q}]=2\left(\begin{array}[]{c|c}BC&0\\ \hline\cr 0&CB\end{array}\right).$
Note that
$BC=\begin{pmatrix}bc&bc^{\prime}\\ b^{\prime}c&b^{\prime}c^{\prime}\end{pmatrix}$
where $B^{T}=(b\;b^{\prime})$ and $C=(c\;c^{\prime})$. But if $BC=0$, then we need either $B=0$ or $C=0$. ∎